OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 2y = 12

OpenStudy (solomonzelman):

set x=0 and solve for y.

OpenStudy (anonymous):

would it be (0, 6) @SolomonZelman ?

OpenStudy (anonymous):

3x - 2y = 12 3(0) - 2y = 12 -2y = 12 y = -12/2 y = -6 solution is : (0,-6)